# New Construction of Locally Perfect Nonlinear Functions with Application to Sequences Sets with Low Ambiguity Zone

### v2

@@ III New Construction of LPNFs
Let $K=p_{1}^{e_{1}}p_{2}^{e_{2}}\cdots p_{s}^{e_{s}}$, where $s\geq 1$, the $p_{i}$ are distinct odd primes and $e_{i}\geq 1$ with $1\leq i\leq s$. Denote $E=\textrm{lcm}\{\varphi(p_{i}^{e_{i}})=p_{i}^{e_{i}-1}(p_{i}-1):1\leq i\leq s\}=NR$ with $N\geq 2$, $N,R\in\mathbb{Z}$. There is an element $\alpha$ of order $N$ in $\mathbb{Z}_{K}^{\ast}$.
@@
For every $i$, $1\leq i\leq s$, The group $\mathbb{Z}_{p_{i}^{e_{i}}}^{\ast}$ is cyclic of order $\varphi(p_{i}^{e_{i}})$, so that $\mathbb{Z}_{p_{i}^{e_{i}}}^{\ast}=\langle\pi_{i}\rangle$ with $\pi_{i}$ as a generator. On the other hand, $\pi=(\pi_{1},\pi_{2},\cdots,\pi_{s})$ is an element of $\mathbb{Z}_{K}^{\ast}$ with order $E=\textrm{lcm}\{\varphi(p_{i}^{e_{i}}):1\leq i\leq s\}$, that is, there exist a cyclic subgroup $\langle\pi\rangle=\{1,\pi,\cdots,\pi^{E-1}\}$ of $\mathbb{Z}_{K}^{\ast}$ with order $E$. Next we write $E=NR$ and put $\alpha=\pi^{R}$, which is an element of $\mathbb{Z}_{K}^{\ast}$ of order $N$. ∎
Construction 1.
Let $K=p_{1}^{e_{1}}p_{2}^{e_{2}}\cdots p_{s}^{e_{s}}$, where $s\geq 1$, the $p_{i}$ are distinct odd primes and $e_{i}\geq 1$ with $1\leq i\leq s$. Denote $E=\textrm{lcm}\{\varphi(p_{i}^{e_{i}})=p_{i}^{e_{i}-1}(p_{i}-1):1\leq i\leq s\}=NR$ with $N\geq 2$, $N,R\in\mathbb{Z}$. We will define the function $f:\,\mathbb{Z}_{N}\rightarrow\mathbb{Z}_{K}$ as $f(x)=\alpha^{x}$, where $\alpha$ is an element of order $N$.
@@
Let $f$ be a function defined in Construction 1. Then we have
(i)  the function $f(x)=\alpha^{x}$ is $\langle N,K,Z_{x},Z_{y}\rangle$LPNF, where $Z_{y}=K$ and $Z_{x}=min\{\frac{p_{i}-1}{\gcd(p_{i}-1,R)}:1\leq i\leq s\}$.
(ii)  the function $f(x)=\alpha^{x}$ is $\langle N,K,Z_{x},Z_{y}\rangle$LPNF, where $Z_{x}=N$ and $Z_{y}=p$, $p$ is the smallest prime factor of $K$.
@@
(i)  For $a\in\mathbb{Z}_{N}\setminus\{0\}$ and $b\in\mathbb{Z}_{K}$, consider the solution $x$ in $\mathbb{Z}_{N}$ of equation
$f(x+a)-f(x)=b,$
which is
@@
[…]
@@
$\Longleftrightarrow$ $p_{i}-1\nmid aR\,\,\,\,\,\,(\mathrm{since\,the\,order\,of}\,\pi\,(mod\,p_{i})\,\mathrm{is}\,p_{i}-1)\,\,\,\,\,\,(1\leq i\leq s)$
@@
Therefore, from the equivalence relation we conclude that if $-Z_{x}<a<Z_{x}$ and $a\neq 0$, where $Z_{x}={\text{min}}\{\frac{p_{i}-1}{\gcd(p_{i}-1,R)}:1\leq i\leq s\}$, then $\alpha^{a}-1\in\mathbb{Z}_{K}^{\ast}$ and reformulate the equation (3) as $\alpha^{x}=b(\alpha^{a}-1)$. Hence, if $b(\alpha^{a}-1)\in\langle\alpha\rangle$, the equation (1) must have a unique solution $x\in\mathbb{Z}_{N}$. Conversely, the equation (1) has no solution in $\mathbb{Z}_{N}$.
(ii)  For any integers $-N<a<N$ with $a\neq 0$ and $-p<b<p$, consider the number of solutions to
$f(x+a)-f(x)\equiv b\,(\rm mod\,K)$ (3)
over $x\in\mathbb{Z}_{N}$, where $p$ denotes the smallest prime factor of $K$. Next we distinguish the two cases $b=0$ and $b\neq 0$.
if $b=0$, then we get $\alpha^{x}(\alpha^{a}-1)\equiv 0\,(\rm mod\,K)$. Since $\alpha\in\mathbb{Z}_{K}^{\ast}$, we have $\alpha^{x}\in\mathbb{Z}_{K}^{\ast}$, and thus $\alpha^{a}-1\equiv 0\,(\rm mod\,K)$, i.e., $\alpha^{a}\equiv 1\,(\rm mod\,K)$. Recall that the order of $\alpha$ mod $K$ is $N$, so $N\,|\,a$, which contradicts to $-N<a<N$ with $a\neq 0$. Therefore, the equation (3) has no solution in $\mathbb{Z}_{N}$.
if $b\neq 0$, then we have $\alpha^{x}(\alpha^{a}-1)\equiv b\,(\rm mod\,K)$. Because $p$ is the smallest prime factor of $K$, for $1\leq i\leq s$, all $p_{i}$ hold $p_{i}\geq p$ and $-p_{i}<b<p_{i}$. By combining $b\neq 0$, we get $p_{i}\nmid b$, that is, $b\not\equiv 0\,(\rm mod\,p_{i})$, which yields $b\in\mathbb{Z}_{p_{i}^{e_{i}}}^{\ast}$ from $\mathbb{Z}_{p_{i}^{e_{i}}}^{\ast}=\mathbb{Z}_{p_{i}^{e_{i}}}\setminus p_{i}\mathbb{Z}_{p_{i}^{e_{i}}}$.Applying the Chinese remainder theorem, we obtain that $b$ is invertible in $\mathbb{Z}_{K}$, i.e., $b\in\mathbb{Z}_{K}^{\ast}$. Note that $\alpha^{x}\in\mathbb{Z}_{K}^{\ast}$, which leads to $\alpha^{a}-1\equiv\alpha^{-x}b\,(\rm mod\,K)$, and hence, it also implies $\alpha^{a}-1\in\mathbb{Z}_{K}^{\ast}$. Together with $\alpha^{x}=(\alpha^{a}-1)^{-1}b\in\mathbb{Z}_{K}^{\ast}$ and the order of $\alpha$ mod $K$ is $N$, we conclude that equation (3) admits a unique solution $x\in\mathbb{Z}_{N}$.
Consequently, the equation (3) has at most a solution. This completes the proof of Theorem 3. ∎
@@
These indicate that $f(x)=2^{x}$ is a locally perfect nonlinear function, which aligns with our Theorem 3 (i).
@@
$\left\{\begin{array}[]{ll}\{f(x-5)-f(x):0\leq x\leq 5\}=\{1,2,4,8,7,5\}\\ \{f(x-4)-f(x):0\leq x\leq 5\}=\{3,6,3,6,3,6\}\\ \{f(x-3)-f(x):0\leq x\leq 5\}=\{7,5,1,2,4,8\}\\ \{f(x-2)-f(x):0\leq x\leq 5\}=\{6,3,6,3,6,3\}\\ \{f(x-1)-f(x):0\leq x\leq 5\}=\{4,8,7,5,1,2\}\\ \{f(x+1)-f(x):0\leq x\leq 5\}=\{1,2,4,8,7,5\}\\ \{f(x+2)-f(x):0\leq x\leq 5\}=\{3,6,3,6,3,6\}\\ \{f(x+3)-f(x):0\leq x\leq 5\}=\{7,5,1,2,4,8\}\\ \{f(x+4)-f(x):0\leq x\leq 5\}=\{6,3,6,3,6,3\}\\ \{f(x+5)-f(x):0\leq x\leq 5\}=\{4,8,7,5,1,2\}\end{array}\right.$
We find that the solutions of $f(x+4)-f(x)=3$ are $1,3,5$. Moreover, $f(x-4)-f(x)=3$, $f(x-2)-f(x)=3$, and $f(x+2)-f(x)=3$ have more than one solution. These imply that $f(x)=2^{x}$ is a locally perfect nonlinear function if $-6<a<6$, $a\neq 0$ and $-3<b<3$, which aligns with our Theorem 3 (ii).
@@
Let $f$ be a function defined in Construction 2. Then we have $f$ is $\langle N,K,Z_{x},Z_{y}\rangle$-LPNF, where $Z_{x}={\text{min}}\{\frac{p_{i}-1}{\gcd(p_{i}-1,R)}:1\leq i\leq s\}$ and $Z_{y}=K_{1}-K+2$.
@@
$f(x+a)-f(x)=\langle\alpha^{x+a}\rangle_{K}-\langle\alpha^{x}\rangle_{K}\equiv b\,(\rm mod\,K_{1}),$ (4)
which has at most one solution $x\in\mathbb{Z}_{N}$ when $-Z_{x}<a<Z_{x}$ with $a\neq 0$ and $-Z_{y}<b<Z_{y}$.
Since $11\leq f(x),f(x+a)\leq K-$ 1, we must have
$-(K-2)\leq f(x+a)-f(x)\leq K-2.$
Thus, from $-(K_{1}-K+2)<b<K_{1}-K+2$ yields
$-(K_{1}-1)\leq f(x+a)-f(x)-b\leq K_{1}-1.$
Furthermore, the equation (4) implies $K_{1}\,|\,f(x+a)-f(x)-b$, and so, it denotes the conventional arithmetic operations applied to integers $\mathbb{Z}$, i.e.,
@@
$f(x+a)-f(x)=\langle\alpha^{x+a}\rangle_{K}-\langle\alpha^{x}\rangle_{K}\equiv b\,(\rm mod\,K).$
@@
$\alpha^{x+a}-\alpha^{x}\equiv b\,(\rm mod\,K).$ (6)
For $-{\text{min}}\{\frac{p_{i}-1}{\gcd(p_{i}-1,R)}:1\leq i\leq s\}<a<{\text{min}}\{\frac{p_{i}-1}{\gcd(p_{i}-1,R)}:1\leq i\leq s\}$ with $a\neq 0$, applying Theorem 3 we can immediately conclude that $\alpha^{a}-1\in\mathbb{Z}_{K}^{\ast}$. Hence, equation (6) has at most a solution in $\mathbb{Z}_{N}$ if $-{\text{min}}\{\frac{p_{i}-1}{\gcd(p_{i}-1,R)}:1\leq i\leq s\}<a<{\text{min}}\{\frac{p_{i}-1}{\gcd(p_{i}-1,R)}:1\leq i\leq s\}$ and $-(K_{1}-K+1)<b<K_{1}-K+1$.
@@
$\left\{\begin{array}[]{ll}\{f(x-5)-f(x):0\leq x\leq 5\}=\{4,9,2,6,1,8\}\\ \{f(x-4)-f(x):0\leq x\leq 5\}=\{3,1,8,7,9,2\}\\ \{f(x-3)-f(x):0\leq x\leq 5\}=\{5,7,9,5,3,1\}\\ \{f(x-2)-f(x):0\leq x\leq 5\}=\{1,8,7,9,2,3\}\\ \{f(x-1)-f(x):0\leq x\leq 5\}=\{2,6,1,8,4,9\}\\ \{f(x+1)-f(x):0\leq x\leq 5\}=\{4,9,2,6,1,8\}\\ \{f(x+2)-f(x):0\leq x\leq 5\}=\{3,1,8,7,9,2\}\\ \{f(x+3)-f(x):0\leq x\leq 5\}=\{5,7,9,5,3,1\}\\ \{f(x+4)-f(x):0\leq x\leq 5\}=\{1,8,7,9,2,3\}\\ \{f(x+5)-f(x):0\leq x\leq 5\}=\{2,6,1,8,4,9\}\end{array}\right.$
These show that the solutions of $f(x+3)-f(x)=5$ and $f(x-3)-f(x)=5$ are $0,3$. we conclude that $f(x)=2^{x}$ is a locally perfect nonlinear function if $-6<a<6$, $a\neq 0$ and $-5<b<5$, which aligns with our Theorem 4.
IV New LAZ sequence sets
By applying the locally perfect nonlinear functions (LPNFs) in Construction 1 and 2 to the above Theorem 2, we directly get the following LAZ sequence sets.

### v3

@@ III Proposed Constructions of LPNFs
Let $K=p_{1}^{e_{1}}p_{2}^{e_{2}}\cdots p_{s}^{e_{s}}$, where $s\geq 1$, the $p_{i}$ are distinct odd primes and $e_{i}\geq 1$ with $1\leq i\leq s$. Denote $E=\textrm{lcm}\{\varphi(p_{i}^{e_{i}})=p_{i}^{e_{i}-1}(p_{i}-1):1\leq i\leq s\}=NR$ with $N\geq 2$, $N,R\in\mathbb{Z}$. Then there is an element $\alpha$ of order $N$ in $\mathbb{Z}_{K}^{\ast}$.
@@
For every $i$, $1\leq i\leq s$, the group $\mathbb{Z}_{p_{i}^{e_{i}}}^{\ast}$ is cyclic of order $\varphi(p_{i}^{e_{i}})$, so that $\mathbb{Z}_{p_{i}^{e_{i}}}^{\ast}=\langle\pi_{i}\rangle$ with $\pi_{i}$ as a generator. On the other hand, $\pi=(\pi_{1},\pi_{2},\cdots,\pi_{s})$ is an element of $\mathbb{Z}_{K}^{\ast}$ with order $E=\textrm{lcm}\{\varphi(p_{i}^{e_{i}}):1\leq i\leq s\}$. Hence, there exist a cyclic subgroup $\langle\pi\rangle=\{1,\pi,\cdots,\pi^{E-1}\}$ of $\mathbb{Z}_{K}^{\ast}$ with order $E$. Since $E=NR$, $\alpha=\pi^{R}$ is an element of $\mathbb{Z}_{K}^{\ast}$ of order $N$. This completes the proof.∎
@@
Let $K=p_{1}^{e_{1}}p_{2}^{e_{2}}\cdots p_{s}^{e_{s}}$, where $s\geq 1$, the $p_{i}$ are distinct odd primes and $e_{i}\geq 1$ with $1\leq i\leq s$. Denote $E=\textrm{lcm}\{\varphi(p_{i}^{e_{i}})=p_{i}^{e_{i}-1}(p_{i}-1):1\leq i\leq s\}=NR$ with $N\geq 2$, $N,R\in\mathbb{Z}$. Define the function $f:\mathbb{Z}_{N}\rightarrow\mathbb{Z}_{K}$ by $f(x)=\alpha^{x}$, where $\alpha\in\mathbb{Z}_{K}^{\ast}$ is an element of order $N$. Then we have
(i)  the function $f(x)=\alpha^{x}$ is $\langle N,K,Z_{x},Z_{y}\rangle$-LPNF, where $Z_{x}=\min\{\frac{p_{i}-1}{\gcd(p_{i}-1,R)}:1\leq i\leq s\}$ and $Z_{y}=K$.
(ii)  the function $f(x)=\alpha^{x}$ is $\langle N,K,Z_{x},Z_{y}\rangle$-LPNF, where $Z_{x}=N$ and $Z_{y}=p$, $p$ is the smallest prime factor of $K$.
@@
(i)  For $a\in\mathbb{Z}_{N}\setminus\{0\}$ and $b\in\mathbb{Z}_{K}$, consider a solution $x$ in $\mathbb{Z}_{N}$ of equation
f(x+a)f(x)=b,f(x+a)-f(x)=b,
then we have
@@
$\alpha^{a}-1\in\mathbb{Z}_{K}^{\ast}$ $\Longleftrightarrow$ $\alpha^{a}-1\in\mathbb{Z}_{p_{i}^{e_{i}}}^{\ast}=\langle\pi_{i}\rangle\,\,\,\,\,\,(\mathrm{for\,all}\,i,\,1\leq i\leq s)$ (2)
@@
$\Longleftrightarrow$ $p_{i}-1\nmid aR\,\,\,\,\,\,(\mathrm{since\,the\,order\,of}\,\pi\,(\bmod\,p_{i})\,\mathrm{is}\,(p_{i}-1))\,\,\,\,\,\,(1\leq i\leq s)$
@@
Therefore, from the equivalence relation we conclude that if $Z_{x}<a<Z_{x}$ and $a\neq 0$, where $Z_{x}={\text{min}}\{\frac{p_{i}-1}{\gcd(p_{i}-1,R)}:1\leq i\leq s\}$, then $\alpha^{a}-1\in\mathbb{Z}_{K}^{\ast}$ and (1) can be transformed into $\alpha^{x}=b(\alpha^{a}-1)^{-1}$. Hence, if $b(\alpha^{a}-1)^{-1}\in\langle\alpha\rangle$, where $\langle\alpha\rangle=\{1,\alpha,\cdots,\alpha^{N-1}\}$ is a cyclic subgroup of $\mathbb{Z}_{K}^{\ast}$ of order $N$, then (1) must have a unique solution $x\in\mathbb{Z}_{N}$. If $b(\alpha^{a}-1)^{-1}\notin\langle\alpha\rangle$, then (1) has no solution in $\mathbb{Z}_{N}$.
(ii)  For any integers N<a<N-N<a<N with $a\neq 0$ and p<b<p-p<b<p, consider the number of solutions to
$f(x+a)-f(x)\equiv b\,(\bmod\,K)$ (3)
over $x\in\mathbb{Z}_{N}$, where $p$ denotes the smallest prime factor of $K$. Next, we distinguish the two cases $b=0$ and $b\neq 0$.
If $b=0$, then we get $\alpha^{x}(\alpha^{a}-1)\equiv 0\,(\bmod\,K)$. Since $\alpha\in\mathbb{Z}_{K}^{\ast}$, we have $\alpha^{x}\in\mathbb{Z}_{K}^{\ast}$, and thus $\alpha^{a}-1\equiv 0\,(\bmod\,K)$, i.e., $\alpha^{a}\equiv 1\,(\bmod\,K)$. Recall that the order of $\alpha$ mod $K$ is $N$, so N|aN\,|\,a, which contradicts to N<a<N-N<a<N with $a\neq 0$. Therefore, the equation (3) has no solution in $\mathbb{Z}_{N}$.
If $b\neq 0$, then we have $\alpha^{x}(\alpha^{a}-1)\equiv b\,(\bmod\,K)$. Because $p$ is the smallest prime factor of $K$, for $1\leq i\leq s$, all $p_{i}$ hold $p_{i}\geq p$ and $p_{i}<b<p_{i}$. By combining $b\neq 0$, we get $p_{i}\nmid b$, that is, $b\not\equiv 0\,(\bmod\,p_{i})$, which yields $b\in\mathbb{Z}_{p_{i}^{e_{i}}}^{\ast}$ from $\mathbb{Z}_{p_{i}^{e_{i}}}^{\ast}=\mathbb{Z}_{p_{i}^{e_{i}}}\setminus p_{i}\mathbb{Z}_{p_{i}^{e_{i}}}$. Applying the Chinese remainder theorem, we obtain that $b$ is invertible in $\mathbb{Z}_{K}$, i.e., $b\in\mathbb{Z}_{K}^{\ast}$. Note that $\alpha^{x}\in\mathbb{Z}_{K}^{\ast}$, which leads to $\alpha^{a}-1\equiv\alpha^{-x}b\,(\bmod\,K)$, and hence, it also implies $\alpha^{a}-1\in\mathbb{Z}_{K}^{\ast}$. Together with $\alpha^{x}=b(\alpha^{a}-1)^{-1}\in\mathbb{Z}_{K}^{\ast}$ and the order of $\alpha$ mod $K$ is $N$, we conclude that equation (3) admits a unique solution $x\in\mathbb{Z}_{N}$.
Consequently, the equation (3) has at most one solution. This completes the proof of Theorem 1. ∎
@@
These indicate that $f(x)=2^{x}$ is a $\langle 20,25,4,20\rangle$-LPNF, which aligns with our Theorem 1 (i).
@@
$\left\{\begin{array}[]{ll}\{f(x-5)-f(x):0\leq x\leq 5\}=\{f(x+1)-f(x):0\leq x\leq 5\}=\{1,2,4,8,7,5\}\\ \{f(x-4)-f(x):0\leq x\leq 5\}=\{f(x+2)-f(x):0\leq x\leq 5\}=\{3,6,3,6,3,6\}\\ \{f(x-3)-f(x):0\leq x\leq 5\}=\{f(x+3)-f(x):0\leq x\leq 5\}=\{7,5,1,2,4,8\}\\ \{f(x-2)-f(x):0\leq x\leq 5\}=\{f(x+4)-f(x):0\leq x\leq 5\}=\{6,3,6,3,6,3\}\\ \{f(x-1)-f(x):0\leq x\leq 5\}=\{f(x+5)-f(x):0\leq x\leq 5\}=\{4,8,7,5,1,2\}\end{array}\right$.
We find that the solutions of f(x+4)f(x)=3f(x+4)-f(x)=3 are $1,3,5$. Moreover, f(x4)f(x)=3f(x-4)-f(x)=3, f(x2)f(x)=3f(x-2)-f(x)=3, and f(x+2)f(x)=3f(x+2)-f(x)=3 have more than one solution. These imply that $f(x)=2^{x}$ is a $\langle 6,9,6,3\rangle$-LPNF, which aligns with our Theorem 1 (ii).
@@
Let $f$ be a function defined in Construction 1. Then we have $f$ is an $\langle N,K_{1},Z_{x},Z_{y}\rangle$-LPNF, where $Z_{x}={\text{min}}\{\frac{p_{i}-1}{\gcd(p_{i}-1,R)}:1\leq i\leq s\}$ and $Z_{y}=K_{1}-K+2$.
@@
$f(x+a)-f(x)=\langle\alpha^{x+a}\rangle_{K}-\langle\alpha^{x}\rangle_{K}\equiv b\,(\bmod\,K),$ (4)
which has at most one solution $x\in\mathbb{Z}_{N}$ when $Z_{x}<a<Z_{x}$ with $a\neq 0$ and $Z_{y}<b<Z_{y}$.
Since $1\leq f(x),f(x+a)\leq K-1$, we must have
$(K-2)\leq f(x+a)-f(x)\leq K-2.$
Thus, from $(K_{1}-K+2)<b<K_{1}-K+2$ yields
$(K_{1}-1)\leq f(x+a)-f(x)-b\leq K_{1}-1.$
Furthermore, the equation (4) implies $K_{1}\,|\,(f(x+a)-f(x)-b)$, and so, it denotes the conventional arithmetic operations applied to integers $\mathbb{Z}$, i.e.,
@@
$f(x+a)-f(x)=\langle\alpha^{x+a}\rangle_{K}-\langle\alpha^{x}\rangle_{K}\equiv b\,(\bmod\,K).$
@@
$\alpha^{x+a}-\alpha^{x}\equiv b\,(\bmod\,K).$ (6)
For ${\min}\{\frac{p_{i}-1}{\gcd(p_{i}-1,R)}:1\leq i\leq s\}<a<{\min}\{\frac{p_{i}-1}{\gcd(p_{i}-1,R)}:1\leq i\leq s\}$ with $a\neq 0$, applying Theorem 1, we can immediately conclude that $\alpha^{a}-1\in\mathbb{Z}_{K}^{\ast}$. Hence, equation (6) has at most a solution in $\mathbb{Z}_{N}$ if ${\min}\{\frac{p_{i}-1}{\gcd(p_{i}-1,R)}:1\leq i\leq s\}<a<{\min}\{\frac{p_{i}-1}{\gcd(p_{i}-1,R)}:1\leq i\leq s\}$ and $(K_{1}-K+2)<b<(K_{1}-K+2)$.
@@
$\left\{\begin{array}[]{ll}\{f(x-5)-f(x):0\leq x\leq 5\}=\{f(x+1)-f(x):0\leq x\leq 5\}=\{4,9,2,6,1,8\}\\ \{f(x-4)-f(x):0\leq x\leq 5\}=\{f(x+2)-f(x):0\leq x\leq 5\}=\{3,1,8,7,9,2\}\\ \{f(x-3)-f(x):0\leq x\leq 5\}=\{f(x+3)-f(x):0\leq x\leq 5\}=\{5,7,9,5,3,1\}\\ \{f(x-2)-f(x):0\leq x\leq 5\}=\{f(x+4)-f(x):0\leq x\leq 5\}=\{1,8,7,9,2,3\}\\ \{f(x-1)-f(x):0\leq x\leq 5\}=\{f(x+5)-f(x):0\leq x\leq 5\}=\{2,6,1,8,4,9\}\end{array}\right$.
These show that the solutions of f(x+3)f(x)=5f(x+3)-f(x)=5 and f(x3)f(x)=5f(x-3)-f(x)=5 are $0,3$. we conclude that $f(x)=2^{x}$ is a $\langle 6,10,6,5\rangle$-LPNF, which aligns with our Theorem 2.
IV Proposed LAZ sequence sets
By applying the LPNFs in Theorems 1 and 2 to the above Lemma 2, we directly get the following LAZ sequence sets. Meanwhile, we also analyze optimality of these LAZ sequence sets.